# Cosmological Stasis from Field-Dependent Decay

### v1

@@ 4 Dynamical properties of the system
Notice that both fixed points can satisfy the physical condition $10<\Omega_{\chi}<$ 1. For certain combinations of $\{x,f,\lambda^{\prime}\}$, the two fixed points can even be physical simultaneously.
@@
Two scans of $\overline{\Omega}_{\chi}$ and the stability of the corresponding fixed point across the $\lambda^{\prime}$-$f$ plane are shown in Fig. 3 for $x=2$ and $x=4$. In both panels, the rainbow-colored region represents the parameter space in which $10<\overline{\Omega}_{\chi}<$ 1, as indicated by the thin black dotted curves. In addition, the thick black curves divide this region into three parts, corresponding to an unstable saddle point, a stable spiral point, and a stable node, respectively.
In both panels, we notice that $f$ is bounded from below, while the lower bound (e.g.$f>1/2$ for $x=2$) increases slowly as $x$ increases. For a fixed value of $f$, the range of $\lambda^{\prime}$ that leads to $10<\overline{\Omega}_{\chi}<$ 1 is finite. However, not all values of $\lambda^{\prime}$ within this range give rise to a stable fixed point. The fixed point tends to be unstable for relatively small $\lambda^{\prime}$. However, as $\lambda^{\prime}$ increases, the fixed point first becomes a stable spiral point and then a stable node if $\lambda^{\prime}$ increases further. Such behavior is consistent with what we have observed in the left panel Fig. 2. In addition, we also observe that the range of $\lambda^{\prime}$ that gives rise to a stable fixed point expands as $f$ increases. However, the range of $\overline{\Omega}_{\chi}$ that this range covers also shrinks as $f$ increases, as we can see from the dotted black contours — the stasis solution becomes increasingly matter-like as $f\gtrsim\mathcal{O}(10)$. Comparing across the two panels of Fig. 3, we also notice that, for a fixed value of $f$, while the range of $\lambda^{\prime}$ that corresponds to a physical $\overline{\Omega}_{\chi}$ shrinks as $x$ increases, the range of $\overline{\Omega}_{\chi}$ that is associated with a stable fixed-point solution expands on the contrary.
@@
To illustrate the global behavior, we fix $\{x,f\}=\{2,1\}$ and present in Fig. 4 cases with different different initial values $\{\Omega_{\chi}^{(0)},\tilde{\Gamma}_{\chi}^{(0)},\Upsilon^{(0)}\}$. In particular, we choose two benchmark values $\overline{\Omega}_{\chi}=0.75$ and $\overline{\Omega}_{\chi}=0.9$ that correspond to a stable spiral point and a stable node, respectively. We then vary the initial value $\Omega_{\chi}^{(0)}$ in the top panels, $\tilde{\Gamma}_{\chi}^{(0)}$ in the middle panels, and $\Upsilon^{(0)}$ in the bottom panels, while holding fixed the other two initial conditions in each row. For both benchmarks, we can see that the evolution converges to the expected stasis values $\{\overline{\Omega}_{\chi},\overline{\tilde{\Gamma}}_{\chi},\overline{\Upsilon}\}$. In particular, all dashed curves in Fig. 4 exhibit intrinsic oscillatory behavior due to the non-vanishing imaginary parts in the eigenvalues of the Jacobian matrix at the fixed point. On the other hand, the solid curves all converge to the stasis value asymptotically after some initial evolution that may temporarily drive the evolution away from stasis. Overall, we find that stasis emerges as a global attractor of the dynamical system as long as the corresponding fixed point is stable, and different choices of initial conditions only affect how the system approaches stasis.
@@ 5.1 Realizing the Hubble mass term
$\Gamma_{\chi\to\phi\phi}$ $=$ $\frac{1}{8192\pi^{5}}\left(\frac{\Lambda_{\rm min}}{\Lambda_{3}}\right)^{2}\left(\frac{m_{\psi}}{m_{\chi}}\right)^{2}m_{\chi}\,,$ (54)
in which $\Lambda_{\rm min}\equiv\min\left\{\Lambda_{1},\Lambda_{2},\Lambda_{3}\right\}$ is the cut-off scale of the fermion loop. Apparently, for $y\ll(\phi/\Lambda_{3})^{2}$, the decay width $\Gamma_{\chi\to\bar{\psi}\psi}$ reduces to the form we propose in Eq. (15), in which we can identify the parameter $C_{\Gamma}=m_{\chi}/(8\pi\Lambda_{3}^{4})$. In addition, to achieve stasis, the field-dependent decay mechanism also requires that $\Gamma_{\chi\to\bar{\psi}\psi\phi\phi}$ and $\Gamma_{\chi\to\phi\phi}$ are subdominant compared to $\Gamma_{\chi\to\bar{\psi}\psi}$. Taken together, these requirements, in turn, lead to conditions on the field value $\phi$ that
@@
$\phi$ $\gg$ $0.056\times\left(\frac{m_{\psi}}{m_{\chi}}\right)^{1/2}\sqrt{\Lambda_{3}\Lambda_{\rm min}}\,.$ (55)
@@ 5.2 Exiting stasis
Figure 6: Time evolution of $\Omega_{\chi}$. The benchmark parameters are chosen such that $m_{\psi}=1$ GeV, $m_{\chi}=10^{10}$ GeV, $m_{\phi}=100$ eV, $\zeta=1/16$, $g=0.01$, $y=4\times 10^{-10}$, $\Lambda_{1}=9.2\times 10^{18}$ Gev, $\Lambda_{2}=1.5\times 10^{15}$ GeV , $\Lambda_{3}=10^{16}$ GeV, $H^{(0)}=250$ GeV, $\rho_{\gamma}^{(0)}=0$, $\phi(t^{(0)})=10^{14}$ GeV and $\dot{\phi}(t^{(0)})=0$. The above parameter choices correspond to $f=1$ and yield a stasis solution with $\overline{\Omega}_{\chi}=0.9$ as indicated by the horizontal dotted line. The three different solid curves represent the different behaviors if we turn on or off $y$ or $m_{\phi}$. The black, blue, purple, and red vertical lines indicate the critical instants $t^{(st)}$, $t_{\mathrm{osc}}$, $t_{y}$, and $t_{\mathrm{RH}}$, respectively.
@@ 5.3 Duration of stasis
$\approx$ $\begin{cases}\displaystyle\frac{\overline{\kappa}}{3}\left[\log(4\pi)+4\log\left(\frac{\Lambda_{3}}{\phi(t^{(\mathrm{st})})}\right)-\log\left(m_{\chi}t^{(\mathrm{st})}\right)\right]&\text{ for Case I}\,,\\ \displaystyle\frac{2}{3}\left[\log(4\pi)+\log\left(\frac{m_{\phi}}{m_{\chi}}\right)-2\log(y)\right]&\text{ for Case II}\,,\end{cases}$
@@
where the superscript “--” or “$+$” denotes the instant immediately before or after the reheating time $t_{\rm RH}$. Together with Eqs. (66) and (67), the energy density of the universe at reheating can be estimated from the matter and radiation energy densities at $t^{(\mathrm{st})}$ by scaling down to $t_{\mathrm{RH}}^{-}$, which leads to
@@ 5.4 Constraints from the relic abundance of $\phi$
    $t_{y}<t_{\mathrm{RH}}<t_{\mathrm{MRE}}<t_{\mathrm{osc}}$: In this case, $\phi$ continues to slow roll even after the $t_{\rm MRE}$. Such a case only happens when the potential of $\phi$ continues to be dominated by the Hubble mass term after MRE. However, since the Hubble mass term originates from the coupling of $\phi$ to the energy densities of different components, whereas $\chi$ has already decayed after $t_{\rm RH}$, the only source that can contribute to the Hubble mass term after reheating is the coupling to the thermal bath through $\psi$. On the other hand, the validity condition for the Hubble mass term, i.e.$m_{\psi,\mathrm{th}}^{2}=\zeta g^{2}T^{2}\gg m_{\psi}^{2}$, suggests that $\psi$ needs to stay relativistic and in thermal equilibrium with the SM thermal bath after $t_{\rm MRE}$. Such a situation will severely violates existing bounds on the effective number of neutrino species $N_{\rm eff}$ (see, e.g., Ref. Planck:2018vyg ). We shall, therefore, exclude this case. Note that, since $t_{\rm osc}\sim 1/m_{\phi}$, while $T_{\rm MRE}\sim 1~{}{\rm eV}$, the constraint which ensures that this case will never occur can be roughly given by $m_{\phi}\gtrsim 10^{-28}~{}{\rm eV}$.
@@ 5.5 Summary of constraints
Figure 7: Parameter scan in the $m_{\phi}$-$y$ plane. In both panels, the color coding indicates the duration of stasis $\mathcal{N}_{s}$, while the dashed black line separates the region for which the universe enters an RD or MD epoch after stasis ends. The white contours represent the reheating temperature $T_{\rm RH}$, and the yellow contour indicates the fraction of dark matter in the form of the oscillating zero-momentum mode of $\phi$. Moreover, the blue and red dashed lines correspond to the values of $y$ below which $\Gamma_{\chi\to\bar{\psi}\psi\phi\phi}$ and $\Gamma_{\chi\to\phi\phi}$ become larger than $\Gamma_{\chi\to\bar{\psi}\psi}^{\rm min}$. Finally, the purple line represents the value of $y$ below which the fraction of energy density in radiation carried by $\phi$ particles with non-zero momenta at reheating, i.e.$\left.{\rho_{\phi_{k}}}/{\rho_{\gamma}}\right|_{t_{\mathrm{RH}}^{+}}$, exceeds $1\%$. In the left panel, the values of the parameters are the same as the ones chosen in Fig. 6 except for $m_{\phi}$ and $y$, which we allow to vary. The benchmark in Fig. 6 is represented by the red star. The right panel is similar to the left panel except that $m_{\chi}=10^{12}$ GeV and $H^{(0)}=25\times 10^{3}$ GeV. In both panels, we have $\overline{\Omega}_{\chi}=0.9$.
@@ 6 Conclusion
Cosmological stasis has been shown to be a possible prediction of many BSM models within a wide range of parameter space. In this paper, we have demonstrated that a stasis between matter and radiation in the early universe can also arise when the decay of the matter component is regulated by the dynamics of a scalar field $\phi$ with negligible abundance. Unlike previous literature, such a stasis does not require the existence of a large tower with an extended mass spectrum, nor does it rely on the thermalization of the matter fields. Instead, the matter/radiation stasis in our work arises due to the interplay between the field-dependent coupling, which controls the decay width, and the potential of $\phi$, which depends on the Hubble rate. Through stability analysis, we have found that, depending on the value of $\overline{\Omega}_{\chi}$, the associated fixed point is either a stable global attractor or an unstable solution that exhibits undamped oscillatory behavior. In the stable regime, we showed that the system can approach stasis either asymptotically or exhibit intrinsic underdamped oscillations. Such properties are quite distinct from all previous stasis scenarios in which stasis is typically a stable node for the entire physical range $10<\overline{\Omega}_{\chi}<$ 1, and the oscillatory behavior observed in the tower-based stasis scenarios Dienes:2023ziv can only appear when the mass spectrum is no longer effectively continuous for relevant timescales.
@@
Aside from the constraints needed for joining the standard $\Lambda$CDM timeline, we have also analyzed the conditions required for the self-consistency of our stasis scenario. One key ingredient of our stasis mechanism is that the decay width $\Gamma_{\chi}$ is dominated by the field-dependent decay channel $\chi\to\bar{\psi}\psi$ for which the field-dependent coupling $(\phi/\Lambda_{3})^{2}$ must also be dominant. In addition, scattering processes involving $\chi$ and $\phi$ must also be suppressed to avoid thermalization. Another crucial ingredient is the validity of the Hubble mass term in the scalar potential $V(\phi)$. This validity is ensured if the thermal mass $m_{\psi,{\rm th}}$ dominates over the classical mass $m_{\psi}$ during stasis, which also ensures that the $\psi$ particles are relativistic. As we have seen in Fig. 7, all the phenomenological and model-consistency constraints can be satisfied in the viable region of the parameter space.
@@
The existence of such a stasis epoch may lead to interesting phenomenological implications, many of which have already been mentioned in Ref. Dienes:2021woi ; Dienes:2022zgd ; Dienes:2023ziv ; Dienes:2024wnu ; Batell:2024dsi . In our explicit model, since $\phi$ is a potential dark-matter candidate, its spectrum of density perturbations today may contain imprints of the earlier non-standard expansion history. It is, therefore, of great interest to study how the growth of different perturbation modes are affected by the stasis epoch and whether or not the perturbation spectrum in our scenario may contain distinguishable features that differ from those studied in other stasis scenarios stasispert .
@@ Acknowledgements.
We thank Michael Ratz for initial collaboration and insightful discussions. We also thank Keith R. Dienes, Max Fieg, Minyuan Jiang, and Brooks Thomas for fruitful discussions. FH would also like to acknowledge the hospitality of the Center for High Energy Physics, Peking University. The research activities of FH are supported in part by ISF Grant 1784/20 and by MINERVA Grant 714123. The work of V.K.-P. was supported by the U.S. National Science Foundation under Grant No. PHY-2210283 and by UC-MEXUS-CONACyT grant No. CN-20-38.

### v2

@@ 4 Dynamical properties of the system
Notice that both fixed points can satisfy the physical condition $0<\Omega_{\chi}<1$. For certain combinations of $\{x,f,\lambda^{\prime}\}$, the two fixed points can even be physical simultaneously.
@@
Two scans of $\overline{\Omega}_{\chi}$ and the stability of the corresponding fixed point across the $\lambda^{\prime}$-$f$ plane are shown in Fig. 3 for $x=2$ and $x=4$. In both panels, the rainbow-colored region represents the parameter space in which $0<\overline{\Omega}_{\chi}<1$, as indicated by the thin black dotted curves. In addition, the thick black curves divide this region into three parts, corresponding to an unstable saddle point, a stable spiral point, and a stable node, respectively.
In both panels, we notice that $f$ is bounded from below, while the lower bound (e.g.f>1/2f>1/2 for $x=2$) increases slowly as $x$ increases. For a fixed value of $f$, the range of $\lambda^{\prime}$ that leads to $0<\overline{\Omega}_{\chi}<1$ is finite. However, not all values of $\lambda^{\prime}$ within this range give rise to a stable fixed point. The fixed point tends to be unstable for relatively small $\lambda^{\prime}$. However, as $\lambda^{\prime}$ increases, the fixed point first becomes a stable spiral point and then a stable node if $\lambda^{\prime}$ increases further. Such behavior is consistent with what we have observed in the left panel Fig. 2. In addition, we also observe that the range of $\lambda^{\prime}$ that gives rise to a stable fixed point expands as $f$ increases. However, the range of $\overline{\Omega}_{\chi}$ that this range covers also shrinks as $f$ increases, as we can see from the dotted black contours — the stasis solution becomes increasingly matter-like as $f\gtrsim\mathcal{O}(10)$. Comparing across the two panels of Fig. 3, we also notice that, for a fixed value of $f$, while the range of $\lambda^{\prime}$ that corresponds to a physical $\overline{\Omega}_{\chi}$ shrinks as $x$ increases, the range of $\overline{\Omega}_{\chi}$ that is associated with a stable fixed-point solution expands on the contrary.
@@
To illustrate the global behavior, we fix $\{x,f\}=\{2,1\}$ and present in Fig. 4 cases with different initial values $\{\Omega_{\chi}^{(0)},\tilde{\Gamma}_{\chi}^{(0)},\Upsilon^{(0)}\}$. In particular, we choose two benchmark values $\overline{\Omega}_{\chi}=0.75$ and $\overline{\Omega}_{\chi}=0.9$ that correspond to a stable spiral point and a stable node, respectively. We then vary the initial value $\Omega_{\chi}^{(0)}$ in the top panels, $\tilde{\Gamma}_{\chi}^{(0)}$ in the middle panels, and $\Upsilon^{(0)}$ in the bottom panels, while holding fixed the other two initial conditions in each row. For both benchmarks, we can see that the evolution converges to the expected stasis values $\{\overline{\Omega}_{\chi},\overline{\tilde{\Gamma}}_{\chi},\overline{\Upsilon}\}$. In particular, all dashed curves in Fig. 4 exhibit intrinsic oscillatory behavior due to the non-vanishing imaginary parts in the eigenvalues of the Jacobian matrix at the fixed point. On the other hand, the solid curves all converge to the stasis value asymptotically after some initial evolution that may temporarily drive the evolution away from stasis. Overall, we find that stasis emerges as a global attractor of the dynamical system as long as the corresponding fixed point is stable, and different choices of initial conditions only affect how the system approaches stasis.
@@ 5.1 Realizing the Hubble mass term
$\Gamma_{\chi\to\phi\phi}$ $=$ $\frac{1}{8192\pi^{5}}\left(\frac{\Lambda_{\rm min}}{\Lambda_{3}}\right)^{2}\left(\frac{m_{\psi}+m_{\psi,\rm th}}{m_{\chi}}\right)^{2}m_{\chi}\,$, (53)
in which $\Lambda_{\rm min}\equiv\min\left\{\Lambda_{1},\Lambda_{2},\Lambda_{3}\right\}$ is the cut-off scale of the fermion loop. Note that, at one loop, the Yukawa term and its field-dependent counterpart give rise to a thermal correction $m_{\chi,\rm th}^{2}\approx\frac{1}{4}\left[y+(\phi/\Lambda_{3})^{2}\right]^{2}T^{2}$ to the mass of $\chi$ particles (see e.g., the last line of Eq. (28) in Ref. Carrington:1991hz ). The consistency of our scenario therefore requires that $m_{\chi}\gg m_{\chi,\rm th}$. To be conservative, we shall require that this condition is satisfied even for the highest temperature $T_{\rm max}$ attained in our numerical results. We emphasize that, since the temperature $T$ decreases with time monotonically except for the transient epoch before reaching $T_{\rm max}$, a violation of this consistency constraint does not necessarily prevent stasis from occurring. Instead, it may only abridge the duration of stasis. Nevertheless, we have checked all our numerical results such that this consistency constraint is always satisfied. We therefore do not include $m_{\chi,\rm th}$ in Eqs. (51), (52) and (53). On the other hand, the mass of $\psi$ particles in Eq. (53) does contain its thermal corrections since we have assumed that the thermal mass of $\psi$ is dominant. As a sanity check, we find that $m_{\chi}\gg T$ at all times in all our cases. Thus, the decay widths in the equations above are always kinematically accessible.
Apparently, for $y\ll(\phi/\Lambda_{3})^{2}$, the decay width $\Gamma_{\chi\to\bar{\psi}\psi}$ reduces to the form we propose in Eq. (15), in which we can identify the parameter $C_{\Gamma}=m_{\chi}/(8\pi\Lambda_{3}^{4})$. In addition, to achieve stasis, the field-dependent decay mechanism also requires that $\Gamma_{\chi\to\bar{\psi}\psi\phi\phi}$ and $\Gamma_{\chi\to\phi\phi}$ are subdominant compared to $\Gamma_{\chi\to\bar{\psi}\psi}$. Taken together, these requirements, in turn, lead to conditions on the field value $\phi$ that
@@
$\phi$ $\gg$ $0.056\times\left(\frac{m_{\psi}+m_{\psi,\rm th}}{m_{\chi}}\right)^{1/2}\sqrt{\Lambda_{3}\Lambda_{\rm min}}\,.$ (54)
Note that, due to the thermal correction $m_{\psi,\rm th}\sim gT$, the third line of the above inequality is time-dependent. However, while $\phi\sim t^{-1/4}$ during stasis, the thermal correction $m_{\psi,\rm th}\sim T\sim\rho_{\gamma}^{1/4}\sim\overline{\Omega}_{\gamma}^{1/4}H^{1/2}\sim t^{-1/2}$. As a result, in the limit $m_{\psi,\rm th}\gg m_{\psi}$, both sides of the corresponding inequality scale with time in the same manner. Indeed, in this limit, both $\Gamma_{\chi\to\psi\bar{\psi}}$ and $\Gamma_{\chi\to\phi\phi}$ scale like $H$ during stasis . Therefore, when the thermal contributions dominate the mass of $\psi$, to ensure the validity of this inequality, it is sufficient to require that $\Gamma_{\chi\to\bar{\psi}\psi}\gg\Gamma_{\chi\to\phi\phi}$ at the maximum temperature during stasis. We have verified that this condition holds in all numerical results presented in this work. However, as we shall show in Sec. 5.3, the tree-level contribution to $\Gamma_{\chi\to\bar{\phi}\phi}$ can impact the duration of stasis since $\Gamma_{\chi\to\bar{\psi}\psi}$ decreases with time and can become comparable to $\Gamma_{\chi\to\phi\phi}$ at later times.
@@ 5.2 Exiting stasis
Figure 6: Time evolution of $\Omega_{\chi}$. The benchmark parameters are chosen such that $m_{\psi}=1\,{\rm GeV}$, $m_{\chi}=10^{10}\,{\rm GeV}$, $m_{\phi}=100\,{\rm eV}$, $\zeta=1/16$, $g=0.01$, $y=4\times 10^{-10}$, $\Lambda_{1}=9.2\times 10^{18}\,{\rm GeV}$, $\Lambda_{2}=1.5\times 10^{15}\,{\rm GeV}$, $\Lambda_{3}=10^{16}\,{\rm GeV}$, $H^{(0)}=250\,{\rm GeV}$, $\rho_{\gamma}^{(0)}=0$, $\phi(t^{(0)})=10^{14}\,{\rm GeV}$ and $\dot{\phi}(t^{(0)})=0$. The above parameter choices correspond to $f=1$ and yield a stasis solution with $\overline{\Omega}_{\chi}=0.9$ as indicated by the horizontal dotted line. The three different solid curves represent the different behaviors if we turn on or off $y$ or $m_{\phi}$. The black, blue, purple, and red vertical lines indicate the critical instants $t^{(\mathrm{st})}$, $t_{\mathrm{osc}}$, $t_{y}$, and $t_{\mathrm{RH}}$, respectively.
@@ 5.3 Duration of stasis
[…]
@@
where the superscript “-” or “++” denotes the instant immediately before or after the reheating time $t_{\rm RH}$. Together with Eqs. (65) and (66), the energy density of the universe at reheating can be estimated from the matter and radiation energy densities at $t^{(\mathrm{st})}$ by scaling down to $t_{\mathrm{RH}}^{-}$, which leads to
@@ 5.4 Constraints from the relic abundance of $\phi$
[…]
@@ 5.5 Summary of constraints
[…]
@@ 6 Conclusion
Cosmological stasis has been shown to be a possible prediction of many BSM models within a wide range of parameter space. In this paper, we have demonstrated that a stasis between matter and radiation in the early universe can also arise when the decay of the matter component is regulated by the dynamics of a scalar field $\phi$ with negligible abundance. Unlike previous literature, such a stasis does not require the existence of a large tower with an extended mass spectrum, nor does it rely on the thermalization of the matter fields. Instead, the matter/radiation stasis in our work arises due to the interplay between the field-dependent coupling, which controls the decay width, and the potential of $\phi$, which depends on the Hubble rate. Through stability analysis, we have found that, depending on the value of $\overline{\Omega}_{\chi}$, the associated fixed point is either a stable global attractor or an unstable solution that exhibits undamped oscillatory behavior. In the stable regime, we showed that the system can approach stasis either asymptotically or exhibit intrinsic underdamped oscillations. Such properties are quite distinct from all previous stasis scenarios in which stasis is typically a stable node for the entire physical range $0<\overline{\Omega}_{\chi}<1$, and the oscillatory behavior observed in the tower-based stasis scenarios Dienes:2023ziv can only appear when the mass spectrum is no longer effectively continuous for relevant timescales.
@@
[…]
@@
The existence of such a stasis epoch may lead to interesting phenomenological implications, many of which have already been mentioned in Ref. Dienes:2021woi ; Dienes:2022zgd ; Dienes:2023ziv ; Dienes:2024wnu ; Batell:2024dsi . In our explicit model, since $\phi$ is a potential dark-matter candidate, its spectrum of density perturbations today may contain imprints of the earlier non-standard expansion history. It is, therefore, of great interest to study how the growth of different perturbation modes are affected by the stasis epoch and whether or not the perturbation spectrum in our scenario may contain distinguishable features that differ from those studied in other stasis scenarios Dienes:2025tox .
@@ Acknowledgements.
We thank Michael Ratz for initial collaboration and insightful discussions. We also thank Keith R. Dienes, Max Fieg, Minyuan Jiang, and Brooks Thomas for fruitful discussions. FH would also like to acknowledge the hospitality of the Center for High Energy Physics, Peking University. The research activities of FH are supported by ISF Grant 1784/20, MINERVA Grant 714123, and MINERVA grant, project 7141230301. The work of V.K.-P. was supported by the U.S. National Science Foundation under Grant No. PHY-2210283 and by UC-MEXUS-CONACyT grant No. CN-20-38.